### v1

# Inflationary power spectrum from the Lanczos algorithm

Abstract
The generalized Lanczos algorithm can provide a universal method for constructing the wave function under the group structure of Hamiltonian. Based on this fact, we obtain an open two-mode squeezed state as the quantum origin for the curvature perturbation. In light of this wave function in the open system, we successfully develop a new method to calculate its corresponding power spectrum by using the Bogoliubov transformation. Unlike traditional approaches, we explicitly retain the Bogoliubov coefficients in terms of the squeezing amplitude $r_{k}$ and the squeezing rotation angle $\phi_{k}$. As a result, the power spectrum of the open two-mode squeezed state will match that of the Bunch-Davies vacuum numerically. Furthermore, the derivation of the open two-mode squeezed state relies on the second kind Meixner polynomial (equivalent to the generalized Lanczos algorithm) and the symmetry of the Hamiltonian. Therefore, our research may offer a new insight into the calculation of the correlation functions through a group-theoretic perspective.
I Introduction
[…]
The quantum origin of curvature perturbation has attracted much attention [13, 14]. It motivates us to use alternative methods from quantum information to evaluate the curvature perturbation. One interesting direction is to implement the concept of complexity to study the curvature perturbations. There are various kinds of complexities in quantum information, such as computational complexity [15], Krylov complexity [46], and etc. The computational complexity in inflation was examined in [17, 18, 19, 20, 21, 22, 23]. Among of these, Ref. [24] has studied the Krylov complexity in cosmology with non-trivial sound speed. However, those literatures studied the complexity for the closed system. But our universe is an open system, therefore, one must use the open system approach to study the Krylov complexity in inflation [25, 26]. In [25, 26], the authors have successfully adopted the two-mode squeezed state to study the Krylov complexity in inflation. As Ref. [47] mentioned, the amplification from the quantum level to the classical level in inflation is a quantum squeezing process. Thus, it is quite natural to use the two-mode squeezed state to evaluate the quantum fluctuations in inflation. The two-mode squeezed state has been broadly applied to the pre-inflationary phase [27, 28, 29, 30, 31, 32].
In this work, we will utilize the Bogoliubov transformation to establish a relationship between the two-mode squeezed state and the Bunch-Davies vacuum (BD). The Bogoliubov coefficients, denoted as $\alpha_{k}$ and $\beta_{k}$, are determined according to [33]. However, the wave function we employ is not the standard pure two-mode squeezed state; instead, we refer to it as the open two-mode squeezed state (OTMSS), which is derived from the second kind of Meixner polynomial (equivalent to the generalized Lanczos algorithm) and the symmetries of the Hamiltonian [26]. Following this, we then calculate the power spectrum of the OTMSS within the framework of the generalized Lanczos coefficient. And we find that the numerical results match those in the BD vacuum.
This work is organized as follows. Sec. II will give some foundations of curvature perturbation under the framework of single-field inflation. Sec. III describes the generalized Lanczos algorithm in inflation. In Sec. IV, we will derive the so-called OTMSS. Sec. V will show the explicit formula of $\alpha_{k}$ and $\beta_{k}$. Then, Sec. VI will obtain the numerics of $\phi_{k}$ and $r_{k}$. Based on this, Sec. VII will derive the power spectrum of OTMSS. Finally, the conclusion and discussion are drawn in Sec. VIII.
II Some foundation of inflationary perturbation
In this work, we will follow Ref. [34] to give some foundation for the cosmological perturbation. First, it is the spatially conformal flat Friedmann-Lemaitre-Robertson-Walker (FLRW) metric,
$ds^{2}=a^{2}(\eta)[-(1+\psi(\eta,\vec{x}))d\eta^{2}+(1-\psi(\eta,\vec{x}))d\vec{x}^{2}],$ (1)
where $\psi(\eta,\vec{x})$ is the perturbation of metric and $a(\eta)$ is the scale factor. Meanwhile, we can also define the perturbation of the inflaton as $\phi(x_{\mu})=\phi_{0}(\eta)+\delta\phi(x_{\mu})$, where $\phi_{0}(\eta)$ is the background field of inflaton and the $\delta\phi(x_{\mu})$ can be dubbed as the perturbation of inflaton. Being armed with these perturbations, we can obtain the generically quadratic action as follows,
[…]
where the Hubble rate $H=\dot{a}/a$ (denoting the derivation with the physical time), the conserved curvature perturbations is $\mathcal{R}=\psi+\frac{H}{\phi_{0}}\delta\phi$ at horizon crossing $(k=aH)$ and the Hubble flow expansion $\epsilon=-\dot{H}/H^{2}$, $z=\sqrt{2\epsilon}a$. Furthermore, we rewrite action (2) in terms of the Mukhanov-Sasaki variable $v=z\mathcal{R}$,
[…]
Applying the Euler-Lagrange equation, we obtain the equation of motion in the momentum space,
$v_{k}^{\prime\prime}+(k^{2}-\frac{z^{\prime\prime}}{z})v_{k}=0.$ (4)
One possible solution of Eq. (4) is the well-known BD vacuum,
$v_{BD}(\eta,\vec{k})=\frac{e^{-ik\eta}}{\sqrt{2k}}(1-\frac{i}{k\eta}).$ (5)
However, defining a privileged quantum vacuum state in curved spacetime is fundamentally problematic, as established by [35]. In this context, the BD vacuum emerges as a particular state that asymptotically approaches the Minkowski vacuum at short distances [36]. The complete vacuum characterization requires consideration of both the Poincaré patch of de Sitter spacetime (PdS) and Hadamard regularity conditions. Following standard cosmological practice, we maintain the BD vacuum assumption under PdS invariance and Hadamard conditions, consistent with foundational work [37, 38, 33].
From an operational perspective, establishing the BD vacuum requires two key elements: $(a)$. The initial conditions for Equation (4) in the sub-horizon regime preceding inflation. $(b)$. The harmonic approximation of (4) as a quantum harmonic oscillator.
The initial condition of Eq.(4) is that it can be approximated as a quantum harmonic oscillator in sub-horizon regime before inflation. These conditions demonstrate that the BD vacuum represents a highly specialized choice for inflationary quantum perturbations. We propose that a more fundamental vacuum state might exist, derivable through Hamiltonian symmetry principles and the Lanczos algorithm. This approach could reveal deeper structure in the quantum initial conditions of inflation.
III Lanczos algorithm in inflation
Recent studies [25, 39, 40, 26, 41] have demonstrated that the early universe behaved as a dissipative system, owing to both the violation of energy conservation [42] and the particle production during preheating [43, 44]. Motivated by these findings, we employ the Lanczos algorithm for open quantum systems to evaluate the quadratic action (3). This approach aligns with the generalized Lanczos algorithm framework developed in [45]. Following our previous conventions [25], we express the generalized Lanczos algorithm as
$\mathcal{L}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1}),$ (6)
where the coefficients $c_{n}$ encode the open quantum system dynamics for a given Hamiltonian, while $b_{n}$ (the Lanczos coefficients) characterize the chaotic properties of the dynamical system. The orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$ spans the Krylov space, where it denotes the space of the two-mode squeezed state, and $\mathcal{L}$ denotes the Liouvillian superoperator – which can be represented in terms of creation and annihilation operators analogous to a Hamiltonian operator.
Following our previous work [25, 39, 40], we identify that:
  • The $c_{n}$ terms in (6) correspond to open system effects
  • The $b_{n}$ terms represent closed system dynamics
This formulation allows us to interpret $\mathcal{L}$ as an effective Hamiltonian that acts on two-mode quantum states, enabling explicit separation between open and closed system contributions. The action of $\mathcal{L}$ on suitable quantum states thus provides a mechanism to:
  1. 1.
    Isolate environmental decoherence effects (open system)
  2. 2.
    Preserve unitary evolution (closed system)
For comprehensive discussions of the Lanczos algorithm implementation and Krylov complexity in this context, we refer readers to [46].
[…]
$\pi(\eta,\vec{x})=\frac{\delta L}{\delta v^{\prime}(\eta,\vec{x})}=v^{\prime}-\frac{a^{\prime}}{a}v.$ (7)
Therefore, we could obtain the Hamiltonian
[…]
Utilizing the Fourier decomposition for $v$ and $\pi$,
$\hat{v}(\eta,\vec{x})=\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{1}{2k}}(\hat{c}_{-\vec{k}}^{\dagger}v_{\vec{k}}^{\ast}e^{-i\vec{k\cdot}\vec{x}}+{\hat{c}_{\vec{k}}}v_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}),$ (9)
$\hat{\pi}(\eta,\vec{x})=i\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{k}{2}}(\hat{c}_{-\vec{k}}^{\dagger}u_{\vec{k}}^{\ast}e^{-i\vec{k\cdot}\vec{x}}-\hat{c}_{\vec{k}}u_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}).$ (10)
where $\hat{c}_{-\vec{k}}^{\dagger}$ and $\hat{c}_{\vec{k}}$ denote the creation and annihilation operators, respectively, $u_{k}$ and $v_{k}$ are the function of conformal time $\eta$. Here, we keep two modes $(k,-k)$ since it is invariant under changing from $k$ to $-k$ in action (3) like in [36]. Therefore, the Hamiltonian (8) will become as follows
$\begin{split}\hat{H}=\int{d^{3}k}\hat{H}_{k}=&\int{d^{3}k}\bigg{[}k\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+k\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\\ &+i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}-i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}\bigg{]}.\end{split}$ (11)
where have chosen appropriate normalizations for $v_{\vec{k}}$ and $u_{\vec{k}}$ and this Hamiltonian is of the Weyl-group structure [57]. Then, we can use this Hamiltonian (11) to act on the two-mode orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$, one can obtain
$\begin{split}\hat{H}|\mathcal{O}_{n})=(2n+1)k|\mathcal{O}_{n})+i\frac{z{}^{\prime}}{z}(n+1)|\mathcal{O}_{n+1})\\ -in\big{(}\frac{z{}^{\prime}}{z}\big{)}|\mathcal{O}_{n-1}).\end{split}$ (12)
where $b_{n}^{2}=n^{2}(\frac{z^{\prime}}{z})^{2}$, $c_{n}=i(2n+1)k$ and the prime symbol indicates differentiation with respect to conformal time. The formula of Lanczos coefficient clearly indicates that the early universe is a maximally chaotic, infinity dynamical system as discovered in [25, 39, 40, 26]. For convenience, we will define
$M_{P}^{2}|1-\mu_{1}^{2}|=\bigg{(}\frac{z^{\prime}}{z}\bigg{)}^{2},~{}~{}M_{P}\mu_{2}=k,$ (13)
where we have followed the notation of [25] and restored the unit of Planck mass, $\mu_{2}$ plays a role of dissipative coefficient and thus $b_{n}^{2}=|1-\mu_{1}^{2}|$ and $c_{n}=i(2n+1)\mu_{2}$. As previously mentioned, the $c_{n}$ part corresponds to the open system part of Hamiltonian (11), and the $b_{n}$ matches the closed system part of Hamiltonian (11). Thus, we could split the Hamiltonian into the open part and closed part as follows,
$\begin{split}&H_{open}=M_{P}\mu_{2}(\hat{c}_{-k}^{\dagger}\hat{c}_{-k}+\hat{c}_{k}\hat{c}_{k}^{\dagger}),\\ &H_{closed}=iM_{P}\sqrt{|1-\mu_{1}^{2}|}(\hat{c}_{k}^{\dagger}\hat{c}_{-k}^{\dagger}-\hat{c}_{k}\hat{c}_{-k}).\end{split}$ (14)
For any single field inflationary models, its Hamiltonian can be divided into two parts as shown in Eq. (14), where the information of various models are encoded in $\mu_{2}$ and $\mu_{1}$.
IV wave function
As Ref. [47] pointed out, the amplification from the quantum level to the classical level in inflation is a quantum squeezing process. Therefore, it is natural to use the two-mode squeezed state to evaluate the quantum fluctuations in inflation. However, in this work, we will implement a more generic function to replace this wave function. First and foremost, the so-called OTMSS via Ref. [25] is
$|\mathcal{O}(\eta))=e^{i\mathcal{L}\eta}|0,0\rangle_{-\vec{k},\vec{k}}=\frac{\rm sechr_{k}(\eta)}{1+\mu_{2}\tanh r_{k}(\eta)}\sum_{n=0}^{\infty}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(-\exp(2i\phi_{k}(\eta))\tanh r_{k}(\eta))^{n}}{(1+\mu_{2}\tanh r_{k}(\eta))^{n}}|n,n\rangle_{-\vec{k},\vec{k}},$ (15)
where $|n,n\rangle_{-\vec{k},\vec{k}}$ is the Krylov basis in this work, $\eta$ is the conformal time, $r_{k}$ and $\phi_{k}$ are two squeezing parameters, $\mu_{2}$ and $|1-\mu_{1}^{2}|$ as defined in Eq. (13). Here, we will introduce more information on wave function (15), in which this wave function will describe the quantum fluctuations in inflation. First, this wave function could represent the dissipative effects of one dynamical system denoted by $\mu_{2}$, and it will nicely recover the two-mode squeezed state in the weak dissipative approximation $(\mu_{2}\ll 1)$,
$|\mathcal{O}(\eta))=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}+\mathcal{O}^{n}(\mu_{2}).$ (16)
Thus, we can see that the pure two-mode squeezed state is a wave function that corresponds to a closed system. In later illustrations, we will confirm this point. Secondly, the derivation of the OTMSS (15) only relies on the second kind of Meixner polynomial [48],
$P_{n+1}(x)=(x-\tilde{c}_{n})P_{n}(x)-b_{n}^{2}P_{n-1}(x),$ (17)
where we define $\tilde{c}_{n}=-ic_{n}$, and $x$ denotew the Hamiltonian. The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$. The polynomial $P_{n}(x)=\det(x-\mathcal{L}_{n})$ corresponds to an open system, and $\mathcal{L}_{n}$ represents the Liouvillian superoperator for the $n$-th quantum state. By introducing the natural orthonormal basis $\{e_{n}\}$. We could represent $|P_{n}(x))=\Bigl{(}\prod_{k=1}^{n}b_{k}\Bigr{)}|\mathcal{O}_{n}),\ \ \mbox{and}\ \ \ |x^{n})=\mathcal{L}^{n}|\mathcal{O}).$ Combining Eq. (17), there is an explicit relation as follows,
$b_{n+1}e_{n+1}+b_{n}e_{n-1}=(x-\tilde{c}_{n})e_{n},$ (18)
where $e_{n}=\mathcal{O}_{n}$ in our case and then it is equivalent to the generalized Lanczos algorithm (6). Through this simple illustration, one could explicitly see that the OTMSS only relies on the second kind of Meixner polynomial, describing the dissipative effects of open system (the early universe). Therefore, we could see that the OTMSS (15) is more generic compared with the pure two-mode squeezed state. Finally, we need to emphasize that the OTMSS is model-independent since the information of various models is encoded in $\mu_{2}$ and $|1-\mu_{1}^{2}|$.
V Bogoliubov transformation
Since the power spectrum of BD vacuum is well consistent with observations [6], we will use the Bogoliubov transformation to construct the relation between the BD vacuum and the vacuum of (15). First, the mode function $v_{z}(\eta,\vec{k})$ parametrizes the squeezed vacuum state in time-evolving spacetime. As for the reference BD vacuum mode, it is denoted by $v_{BD}(\eta,\vec{k})$. Then, we could introduce the non-trivial Bogoliubov transformation to relate $v_{z}(\eta,\vec{k})$ and $v_{BD}(\eta,\vec{k})$ as follows,
$v_{z}(\eta,\vec{k})=\alpha_{k}v_{BD}(\eta,\vec{k})+\beta_{k}v_{BD}^{*}(\eta,\vec{k}),$ (19)
where the coefficient of Bogoliubov transformation satisfying with the canonical Wronskian condition,
$|\alpha_{k}|^{2}-|\beta_{k}|^{2}=1.$ (20)
Meanwhile, the annihilation operator of OTMSS can also be defined by
$\hat{c}_{\vec{k}}=\alpha_{k}\hat{a}_{\vec{k}}+\beta_{k}^{*}\hat{a}_{-\vec{k}}^{\dagger}.$ (21)
Then, the we could use the Bogoliubov coefficient to construct the wave function via BD vacuum as follows,
$|0,0\rangle_{OST}=\frac{1}{\mathcal{N}}\exp\bigg{[}-\int\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\frac{\beta_{k}^{*}}{\alpha_{k}^{*}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\bigg{]}|BD\rangle,$ (22)
where $\mathcal{N}$ represents the normalization factor, and $|0,0\rangle_{OST}$ denotes the vacuum of the OTMSS. The essence of Eq. (22) is that one can construct the wave function in the target spacetime using the Bogoliubov transformation and the BD vacuum. Therefore, the key aspect is determining the formulas for $\alpha_{k}$ and $\beta_{k}$ for our specific case. This is also crucial for understanding the relationship between the modes of the squeezed state and the modes of the BD vacuum, as indicated in Eq. (23). In fact, we can apply the squeezed operator $\mathcal{\hat{S}}_{-k,k}=\exp\left(\frac{r_{k}}{2}e^{-i\phi_{k}}\hat{a}_{k}\hat{a}_{-k}-\frac{r_{k}}{2}e^{i\phi_{k}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$ to the annihilation operator of the BD vacuum to derive the coefficients for the Bogoliubov transformation,
$\mathcal{\hat{S}}^{\dagger}\hat{a}\mathcal{\hat{S}}=\hat{c}_{\vec{k}}=\cosh r_{k}\hat{a}_{\vec{k}}-e^{i\phi_{k}}\sinh r_{k}\hat{a}_{-\vec{k}}^{\dagger},$ (23)
Compared with Eq. (21), we can obtain the formula for $\alpha_{k}$ and $\beta_{k}$
$\alpha_{k}=\cosh r_{k},~{}~{}~{}\beta_{k}=-\exp(-i\phi_{k})\sinh r_{k},$ (24)
where $\phi_{k}$ and $r_{k}$ are two parameters depicting the properties for the OTMSS (15) and the formula of $\alpha_{k}$ and $\beta_{k}$ is consistent with [49, 50]. Next, we will proceed with the evolution of $\phi_{k}$ and $r_{k}$ determined by the Hamiltonian (11).
As we discussed, the OTMSS (15) is more generic compared with a pure two-mode squeezed state, including the dissipative effects of an open system. In this work, we will particularly focus on the super Hubble scale, namely, $k\eta$ will be approaching zero. In light of (13), we could see that $\mu_{2}$ (dissipative coefficient) will also be very tiny in this super Hubble scale, meaning that the universe within the super Hubble scale is a weak dissipative system. Consequently, the pure two-mode squeezed state will be a good approximation, and our derivation for (24) is also consistent with each other.
VI Evolution of $\phi_{k}$ and $r_{k}$
To obtain the power spectrum of the OTMSS (15), one needs to rewrite the Bogoliubov coefficients (24) in terms of $k$ as in Refs. [51, 52, 53], where their formulas has made some approximation. To fully capture the information of (15), we will explicitly numerically give the evolution of $r_{k}$ and $\phi_{k}$ in terms of $k$, where the evolution is completely determined by the Schrödinger equation $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$. The resulting evolution equation of $r_{k}$ and $\phi_{k}$ can be denoted by via [26]
$r_{k}^{\prime}=\frac{-M_{P}|1-\mu_{1}^{2}|\sinh 2r_{k}\cos 2\phi_{k}-\sinh 2r_{k}\cdot\mu_{2}^{\prime}}{\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k}},$ (25)
$\phi_{k}^{\prime}=-M_{P}\mu_{2}+\frac{1}{2}M_{P}|1-\mu_{1}^{2}|^{\frac{1}{2}}\sin 2\phi_{k}\Big{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\coth r_{k}+\mu_{2})\Bigr{]}.$ (26)
Compared to the original evolution equations for $\phi_{k}$ and $r_{k}$, we have neglected the contribution from the potential due to the slow-roll conditions. These calculations can be quite complicated; for more details, the reader can refer to the appendix of [26]. The evolution equations (26) for $r_{k}$ and $\phi_{k}$ describe the dynamics during the inflationary period. The power spectrum was observed at the horizon crossing at a fiducial momentum of $k_{*}=0.05~{}\text{Mpc}^{-1}$. To compare with observations, the data from the CMB covers the range from $10^{-4}~{}\text{Mpc}^{-1}$ to $1~{}\text{Mpc}^{-1}$ [6], making it beneficial to fully expand the power spectrum within this range. At the pivot point, we have $k_{*}=aH$, where the value is derived from today’s measurements, allowing us to easily obtain $\eta=-\frac{1}{k}$ within the CMB observational range [54, 55, 56]. Given that $\eta=-\frac{1}{aH}$, the evolution can be transformed accordingly,
$\begin{split}r_{k}^{\prime}=&\ -\frac{\tanh r_{k}(\mu_{2}^{\prime}+M_{P}|1-\mu_{1}^{2}|\cos 2\phi_{k})}{\tanh r_{k}+\mu_{2}},\\ \phi_{k}^{\prime}=&\ -M_{P}\mu_{2}+\frac{1}{2}\sin 2\phi_{k}\Bigl{(}\frac{M_{P}|1-\mu_{1}^{2}|\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\\ &\ +M_{P}\coth r_{k}+M_{P}\mu_{2}\Bigr{)}.\end{split}$ (27)
Being armed with these two equations, we could give their numeric in terms of comoving momentum $k$.
Figure 1: The numeric of squeezed parameter $r_{k}$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Figure 2: The numeric of squeezed parameter $\phi_{k}$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
From Fig. 1 and Fig. 2, one can see that the evolution of $r_{k}$ and $\phi_{k}$ are almost the constants from $10^{-4}~{}Mpc^{-1}$ to $1~{}Mpc^{-1}$. Even, we can further evaluate the number density of OTMSS in terms of BD vacuum,
$\langle 0,0|\hat{c}^{\dagger}_{\vec{k}}\hat{c}_{-\vec{k}}|0,0\rangle=|\beta_{k}|^{2}=\sinh^{2}r_{k},$ (28)
where $|0,0\rangle$ is the BD vacuum and we have implemented Eq. (21). In light of this equation, we can give the numeric of mean number density under the BD vacuum.
Figure 3: The numeric of squeezed parameter $|\beta_{k}|^{2}$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Fig. 3 clearly indicates that mean number density is almost zero since the particles are generated in preheating process, although the number density of particle is non-vanishing. Being armed with the evolution of $r_{k}$ and $\phi_{k}$ in terms of k, one can investigate the power spectrum.
VII Power spectrum of OTMSS
Since the power spectrum of OTMSS (15) is proportional to $|v_{z}|^{2}$. Then, we can obtain the ratio between the power spectrum of the OTMSS and the BD vacuum in light of (23),
$\gamma_{z}=\frac{\Delta_{\mathcal{R}z}^{2}}{\Delta_{\mathcal{R}}^{2}}=|\alpha-\beta|^{2}=\cosh 2r_{k}+\sinh 2r_{k}\cos\phi_{k},$ (29)
where $\Delta_{\mathcal{R}z}^{2}$ denotes the power spectrum for the OTMSS, $\Delta_{\mathcal{R}}^{2}$ represents the power spectrum of BD vacuum and we have used $\lim_{k\eta\to 0}\frac{v_{BD}^{*}}{v_{BD}}=-1$ (super Hubble scale approximation) via BD vacuum (5). Once having this relation, it is quite straightforwardly showing the numeric of (29) in terms of k.
Figure 4: The numeric of squeezed parameter $\gamma_{z}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Fig. 4 indicates that the ratio (29) is almost one after the horizon crossing, inferring that the power spectrum of two-mode squeezed state will be consistent with the BD vacuum. For completeness, we will use the definition of power to confirm this conclusion. Following the notation of [34], the two-point correlation function can be defined as
$\begin{split}\langle|\hat{v}_{z}|^{2}\rangle=&\ \langle 0,0|\hat{v}_{z}^{\dagger}(\eta,0)\hat{v}_{z}(\eta,0)|0,0\rangle\\ =&\ \int d\ln{k}\frac{k^{3}}{2\pi^{2}}|v_{BD}(\eta,\vec{k})|^{2}\gamma_{z}.\end{split}$ (30)
where $\gamma_{z}$ is the ratio (29). Thereafter, we could define the power spectrum of the mode function $v_{z}$ as follows,
$\begin{split}\Delta_{v_{z}}^{2}(\eta,k)=&\ \frac{k^{3}}{2\pi^{2}}|v_{BD}(\eta,\vec{k})|^{2}\gamma_{z}\\ =&\ \Delta_{v_{BD}}^{2}(\eta,k)\gamma_{z},\end{split}$ (31)
To match the observations, we will use the curvature perturbation to calculate the power spectrum, where the curvature perturbation is $\mathcal{R}=v/z$ with $z=\sqrt{2\epsilon}a$ as shown in Sec. II. Thus, we could obtain the power spectrum of curvature perturbation as follows,
$\Delta_{\mathcal{R}z}^{2}(\eta,k)=\frac{1}{2\epsilon a^{2}}\frac{\Delta_{v_{z}}^{2}(\eta,k)}{M_{P}^{2}}=\frac{1}{2\epsilon a^{2}}\frac{\Delta_{v_{BD}}^{2}\gamma_{z}(\eta,k)}{M_{P}^{2}},$ (32)
As for the power spectrum of BD vacuum, it will predict the $\Delta_{\mathcal{R}}^{2}=A_{s}(\frac{k}{k_{*}})^{n_{s}-1}$ where the observational constraints to the amplitude power spectrum $A_{s}$ and spectral index from [6]
$A_{s}=(2.196\pm 0.060)\times 10^{-9},~{}~{}~{}n_{s}=0.9649\pm 0.0042.$ (33)
Based on these observational constraints, we could give the prediction for the power spectrum of OTMSS in light of (32).
Figure 5: The numeric of squeezed parameter $\Delta_{\mathcal{R}z}^{2}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Fig. 5 clearly shows the power spectrum of the OTMSS is almost the same with BD vacuum due to the ratio relation (29) as shown in Fig. 4, meanwhile our evaluation for the spectral index $n_{s}$ is also consistent with the BD vacuum.
Recalling the calculation of the power spectrum for the squeezed state as outlined in [51, 52, 53], we can parametrize the Bogoliubov coefficients, $\alpha_{k}$ and $\beta_{k}$, in terms of $r_{k}$ and $\phi_{k}$. These coefficients will be expressed as functions of the momentum $k$. However, these procedures overlook a crucial aspect: both $r_{k}$ and $\phi_{k}$ evolve over time as determined by the corresponding Schrödinger equation, given by $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$. Therefore, a more accurate approach would be to take the time evolution of $r_{k}$ and $\phi_{k}$ into consideration. Furthermore, we focus on the OTMSS, as described in (15), which represents the open system (the universe is regarded as an open system). Following this approach, our numerical results for the power spectrum are consistent with the BD vacuum. While our derivation cannot be distinguished from the BD vacuum, the wave function in (15) is more general, being model-independent and relying solely on the second Meixner polynomial and the Lanczos algorithm.
VIII Conclusion and Discussion
Under the single field inflation, we have constructed a framework for calculating the power spectrum of two mode squeezed state, where the key ingredient is the OTMSS (15) and the generalized Lanczos algorithm as shown in Sec. III.
In light of the Lanczos algorithm, the Hamiltonian can be separated into components corresponding to open and closed systems, as shown in Eq. (14). During inflation, an amplification process drives quantum phenomena toward classicalization, making the two-mode squeezed state a natural framework for studying curvature perturbations. The squeezed state is constructed via the action of the squeezed operator $\mathcal{\hat{S}}_{-k,k}=\exp\left(u\hat{a}_{k}\hat{a}_{-k}-\bar{u}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$ on the BD vacuum, where $u=\frac{r_{k}}{2}e^{-i\phi_{k}}$ in our case. Crucially, (14) reveals that $\mathcal{\hat{S}}_{-k,k}$ exclusively encodes the closed system dynamics. In contrast, the OTMSS in (15) incorporates environmental interactions through the dissipative coefficient $\mu_{2}$. The standard squeezed state emerges as the leading-order term in the weak dissipation limit ($\mu_{2}\to 0$). From the Schrödinger equation $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$, we derive the evolution equations for $r_{k}(\eta)$ and $\phi_{k}(\eta)$, whose numerical solutions are plotted in Figs. 2 and 1. The ratio between the OTMSS and BD vacuum, defined via Bogoliubov transformations in Eq. (29). As shown in Fig. 4, $\gamma_{k}$ remains unity throughout inflation, consistent with suppressed particle production (mean particle density $\langle n_{k}\rangle\ll 1$) due to preheating dynamics. Consequently, the power spectrum $\Delta_{\mathcal{R}}^{2}(k)$ computed from the OTMSS in (15) matches the BD vacuum prediction, as demonstrated in Fig. 5.
Key insights:
  • Dissipation ($\mu_{2}$) modifies higher-order correlations in the strong dissipative regime but leaves the leading-order power spectrum intact.
  • Observational indistinguishability from BD vacuum arises from $\gamma_{k}\approx 1$, even with $\mu_{2}\neq 0$.
  • Lanczos-based separation of $\hat{H}$ clarifies the open system’s part and closed system’s part.
The derivation of the BD vacuum in (5) relies on solving the mode equation. In contrast, the OTMSS is constructed through the Hamiltonian formalism and Lanczos algorithm without solving equations of motion, where model-dependent information is encoded in the parameters $\mu_{2}$ and $|1-\mu_{1}^{2}|$. Remarkably, the power spectrum prediction from the OTMSS in (15) matches the BD vacuum result, suggesting that squeezed state construction is a more fundamental approach than solving EOM.
The key ingredient in this construction is the generalized displacement operator [57], defined as
$\mathcal{D}(\xi)=\exp\left(\xi\hat{H}-\bar{\xi}\hat{H}^{\dagger}\right),$ (34)
where $\hat{H}$ exhibits specific group structures (e.g., the Weyl algebra in our case) and the squeezed operator belongs this kind of operator. When $\hat{H}$ belongs to a Lie group, the operator $\mathcal{D}(\xi)$ provides a rigorous framework for generating quantum states, surpassing traditional EOM-based methods in universality.
Extensions and implications:
  • Current results apply to single-field inflation. Generalizations to multifield scenarios [58, 59, 60, 61, 62] and $f(R)$ gravity [63, 64] are natural next steps.
  • The $\mu_{2}$-dependence in OTMSS may encode decoherence effects observable in higher-order correlations.
  • Group-theoretic construction offers a unified approach for quantum fluctuations in curved spacetime.
Acknowledgements.
LHL and KHZ are funded by National Natural Science Foundation of China (NSFC) with grant NO. 12165009, Hunan Natural Science Foundation with grant NO. 2023JJ30487 and NO. 2022JJ40340. HQZ is funded by NSFC with grant NO. 12175008.

### v2

@@
 I introduction
-[…]
-The quantum origin of curvature perturbation has attracted much attention [13, 14].
-It motivates us to use alternative methods from quantum information to evaluate the curvature perturbation.
-One interesting direction is to implement the concept of complexity to study the curvature perturbations.
-There are various kinds of complexities in quantum information, such as computational complexity [15], Krylov complexity [46], and etc.
-The computational complexity in inflation was examined in [17, 18, 19, 20, 21, 22, 23].
-Among of these, Ref.
-[24] has studied the Krylov complexity in cosmology with non-trivial sound speed.
-However, those literatures studied the complexity for the closed system.
-But our universe is an open system, therefore, one must use the open system approach to study the Krylov complexity in inflation [25, 26].
-In [25, 26], the authors have successfully adopted the two-mode squeezed state to study the Krylov complexity in inflation.
-As Ref.
-[47] mentioned, the amplification from the quantum level to the classical level in inflation is a quantum squeezing process.
-Thus, it is quite natural to use the two-mode squeezed state to evaluate the quantum fluctuations in inflation.
-The two-mode squeezed state has been broadly applied to the pre-inflationary phase [27, 28, 29, 30, 31, 32].
-In this work, we will utilize the Bogoliubov transformation to establish a relationship between the two-mode squeezed state and the Bunch-Davies vacuum (BD).
-The Bogoliubov coefficients, denoted as $\alpha_{k}$ and $\beta_{k}$, are determined according to [33].
-However, the wave function we employ is not the standard pure two-mode squeezed state; instead, we refer to it as the open two-mode squeezed state (OTMSS), which is derived from the second kind of Meixner polynomial (equivalent to the generalized Lanczos algorithm) and the symmetries of the Hamiltonian [26].
-Following this, we then calculate the power spectrum of the OTMSS within the framework of the generalized Lanczos coefficient.
-And we find that the numerical results match those in the BD vacuum.
+The inflationary paradigm Guth:1980zm ; Starobinsky:1980te ; Linde:1981mu addresses several fundamental issues in standard Big Bang cosmology, including the flatness problem, the horizon problem, and the magnetic monopole problem.
+Quantum fluctuations in the vacuum state during quasi-de Sitter expansion predict the power spectrum of primordial curvature perturbations Mukhanov:1990me ; Martin:2024qnn , which is nearly scale-invariant as constrained by observations of the cosmic microwave background (CMB) Planck:2018jri .
+The advent of high-precision CMB measurements has initiated an era of precision cosmology.
+Consequently, current and forthcoming observations are capable of probing primordial curvature perturbations beyond the predictions of standard inflationary models.
+CORE:2016ymi ; CMB-S4:2020lpa ; LiteBIRD:2022cnt ; Euclid:2023shr ; Mergulhao:2023ukp ; Antony:2024vrx .
+The quantum origin of curvature perturbations has attracted significant attention Balitsky:2014epa ; Dalianis:2024 .
+This motivates the application of alternative methods from quantum information theory to evaluate curvature perturbations.
+One promising direction involves utilizing the concept of complexity to study these perturbations.
+Several types of complexity exist in quantum information, including computational complexity Papadimitriou:2003 , Krylov complexity Parker:2018yvk , among others.
+Computational complexity in the context of inflation has been investigated in Choudhury:2020hil ; Bhargava:2020fhl ; Bhattacharyya:2020rpy ; Adhikari:2021ked ; Li:2023ekd ; Liu:2021nzx ; Li:2021kfq .
+Among these, Ref.
+Adhikari:2022oxr studied Krylov complexity in cosmology with non-trivial sound speed.
+However, these works primarily address closed quantum systems.
+Since our universe is fundamentally an open system, it is essential to adopt an open quantum systems approach to study Krylov complexity during inflation Li:2024kfm ; Zhai:2024odw .
+In Li:2024kfm ; Zhai:2024odw , the authors successfully employed the two-mode squeezed state formalism to analyze Krylov complexity in inflation.
+As emphasized in Grishchuk:1990bj , the amplification of quantum fluctuations to classical scales during inflation is effectively a quantum squeezing process.
+Therefore, the two-mode squeezed state provides a natural framework for describing quantum fluctuations in this context.
+This approach has also been extensively applied to study the pre-inflationary phase Kowalczyk:2024ech ; Vilenkin:1982wt ; Starobinsky:2001kn ; Gessey-Jones:2021yky ; Handley:2016ods ; Contaldi:2003zv .
+In this work, we employ the Bogoliubov transformation to relate the two-mode squeezed state to the Bunch-Davies (BD) vacuum.
+The Bogoliubov coefficients, $\alpha_{k}$ and $\beta_{k}$, are determined according to Bunch:1978yq .
+The wave function used here is not the conventional pure two-mode squeezed state, but rather the open two-mode squeezed state (OTMSS), constructed via the second kind of Meixner polynomial, which is equivalent to the generalized Lanczos algorithm along with the symmetries of the Hamiltonian Zhai:2024odw .
+Within this framework, we compute the power spectrum of the OTMSS using the generalized Lanczos coefficients.
+Our numerical results show agreement with those obtained in the BD vacuum.
 This work is organized as follows.
@@
 In Sec.
-IV, we will derive the so-called OTMSS.
+IV, we will review the so-called OTMSS.
 Sec.
@@
 VII will derive the power spectrum of OTMSS.
-Finally, the conclusion and discussion are drawn in Sec.
+Finally, the conclusion and discussion will be given in Sec.
 VIII.
 II Some foundation of inflationary perturbation
-In this work, we will follow Ref. [34] to give some foundation for the cosmological perturbation. First, it is the spatially conformal flat Friedmann-Lemaitre-Robertson-Walker (FLRW) metric,
+In this work, we will follow Ref. Baumann:2009ds to give some foundation for the cosmological perturbation. First, it is the spatially conformal flat Friedmann-Lemaitre-Robertson-Walker (FLRW) metric,
 $ds^{2}=a^{2}(\eta)[-(1+\psi(\eta,\vec{x}))d\eta^{2}+(1-\psi(\eta,\vec{x}))d\vec{x}^{2}],$ (1)
-where $\psi(\eta,\vec{x})$ is the perturbation of metric and $a(\eta)$ is the scale factor.
-Meanwhile, we can also define the perturbation of the inflaton as $\phi(x_{\mu})=\phi_{0}(\eta)+\delta\phi(x_{\mu})$, where $\phi_{0}(\eta)$ is the background field of inflaton and the $\delta\phi(x_{\mu})$ can be dubbed as the perturbation of inflaton.
-Being armed with these perturbations, we can obtain the generically quadratic action as follows,
-[…]
-where the Hubble rate $H=\dot{a}/a$ (denoting the derivation with the physical time), the conserved curvature perturbations is $\mathcal{R}=\psi+\frac{H}{\phi_{0}}\delta\phi$ at horizon crossing $(k=aH)$ and the Hubble flow expansion $\epsilon=-\dot{H}/H^{2}$, $z=\sqrt{2\epsilon}a$.
-Furthermore, we rewrite action (2) in terms of the Mukhanov-Sasaki variable $v=z\mathcal{R}$,
-[…]
+where $\psi(\eta,\vec{x})$ denotes the metric perturbation and $a(\eta)$ is the scale factor.
+We also define the perturbation of the inflaton field as $\phi(x_{\mu})=\phi_{0}(\eta)+\delta\phi(x_{\mu})$, where $\phi_{0}(\eta)$ is the homogeneous background inflaton and $\delta\phi(x_{\mu})$ represents the inflaton perturbation.
+Using these definitions, the generic quadratic action can be derived as follows:
+$S=\frac{1}{2}\int dtd^{3}xa^{3}\frac{\dot{\phi}}{H^{2}}\bigl{[}\dot{\mathcal{R}}-\frac{1}{a}(\partial_{i}\mathcal{R})^{2}\bigr{]},$ (2)
+where the Hubble parameter is defined as $H=\dot{a}/a$ (denoting the derivation with the physical time), the conserved curvature perturbations is $\mathcal{R}=\psi+\frac{H}{\phi_{0}}\delta\phi$ at horizon crossing $(k=aH)$ and the slow-roll parameter is $\epsilon=-\dot{H}/H^{2}$, which leads to $z=\sqrt{2\epsilon}a$.
+Furthermore, the action (2) can be rewritten in terms of the Mukhanov-Sasaki variable $v=z\mathcal{R}$,
+$S=\frac{1}{2}\int d\eta d^{3}x[v^{\prime 2}-(\partial_{i}v)^{2}+(\frac{z^{\prime}}{z})^{2}v^{2}-2v^{\prime}v\frac{z^{\prime}}{z}].$ (3)
 Applying the Euler-Lagrange equation, we obtain the equation of motion in the momentum space,
@@
 $v_{BD}(\eta,\vec{k})=\frac{e^{-ik\eta}}{\sqrt{2k}}(1-\frac{i}{k\eta}).$ (5)
-However, defining a privileged quantum vacuum state in curved spacetime is fundamentally problematic, as established by [35].
-In this context, the BD vacuum emerges as a particular state that asymptotically approaches the Minkowski vacuum at short distances [36].
-The complete vacuum characterization requires consideration of both the Poincaré patch of de Sitter spacetime (PdS) and Hadamard regularity conditions.
-Following standard cosmological practice, we maintain the BD vacuum assumption under PdS invariance and Hadamard conditions, consistent with foundational work [37, 38, 33].
+However, the definition of a privileged quantum vacuum state in curved spacetime is fundamentally ambiguous, as emphasized in Unruh:1976db .
+In this setting, the BD vacuum arises as a specific state that approximates the Minkowski vacuum at short distances Agullo:2022ttg .
+A complete characterization of the vacuum requires the framework of the Poincaré patch of de Sitter spacetime (PdS) together with Hadamard regularity conditions.
+In accordance with standard cosmological practice, we adopt the BD vacuum under the assumptions of PdS invariance and Hadamard regularity, following the foundational treatments in Chernikov:1968zm ; Tagirov:1972vv ; Bunch:1978yq .
 From an operational perspective, establishing the BD vacuum requires two key elements: $(a)$.
@@
 The harmonic approximation of (4) as a quantum harmonic oscillator.
-The initial condition of Eq.(4) is that it can be approximated as a quantum harmonic oscillator in sub-horizon regime before inflation.
+The initial condition of Eq.
+(4) is that it can be approximated as a quantum harmonic oscillator in sub-horizon regime before inflation.
 These conditions demonstrate that the BD vacuum represents a highly specialized choice for inflationary quantum perturbations.
@@
 III Lanczos algorithm in inflation
-Recent studies [25, 39, 40, 26, 41] have demonstrated that the early universe behaved as a dissipative system, owing to both the violation of energy conservation [42] and the particle production during preheating [43, 44].
+Recent studies Li:2024kfm ; Li:2024iji ; Li:2024ljz ; Zhai:2024odw ; Shandera:2017qkg have demonstrated that the early universe behaved as a dissipative system, owing to both the violation of energy conservation Cheung:2007st and the particle production during preheating Kofman:1997yn ; Kofman:1994rk .
 Motivated by these findings, we employ the Lanczos algorithm for open quantum systems to evaluate the quadratic action (3).
-This approach aligns with the generalized Lanczos algorithm framework developed in [45].
-Following our previous conventions [25], we express the generalized Lanczos algorithm as
+There are multiple approaches to Krylov complexity in open quantum systems.
+In this work, we employ the generalization of the Lanczos algorithm to the Arnoldi iteration via the Heissenberg form Bhattacharjee:2022lzy ; Bhattacharya:2022gbz ; Liu:2022god ; Nizami:2023dkf ; Nandy:2024evd , as this framework has proven effective for generating the OTMSS in our previous studies Li:2024kfm ; Li:2024iji ; Li:2024ljz ; Zhai:2024odw , particularly building on the methodology established in Bhattacharya:2022gbz .
+We begin by introducing the general Lindblad master equation Lindblad:1975ef :
+$\dot{\rho}=-i[H,\rho]+\sum_{k}[L_{k}\rho L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho\}],$ (6)
+where $\rho$ is the density matrix, $H$ is the Hamiltonian, and $L_{k}^{\dagger}$ denotes the jump operator (which corresponds to the annihilation operator in our context).
+Following Bhattacharjee:2022lzy , we assume the open system dynamics occurs in an infinite temperature environment-modeled by the maximally mixed state.
+This serves as a reasonable approximation since the temperature of the inflationary universe is exceedingly high.
+Consequently, we obtain
+$\sum_{k}[L_{k}\rho_{\infty}L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho_{\infty}\}]=0,$ (7)
+where $\rho_{\infty}$ denotes the infinite temperature-state. Under this assumption, the operator dyamics is governed by
+$\mathcal{\hat{O}}(t)=\exp(i\mathcal{L}t)\mathcal{\hat{O}}$, (8)
+where the Heisenberg-picture Lindbladian $\mathcal{L}$ is given by Bhattacharya:2022gbz ,
+$\mathcal{L}=\mathcal{L}_{H}+\mathcal{L}_{D},~~\mathcal{L}_{H}\mathcal{\hat{O}}=[H,\mathcal{\hat{O}}],~~\mathcal{L}_{D}=\sum_{k}[L_{k}\mathcal{\hat{O}}L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\mathcal{\hat{O}}\}$] (9)
+where $\mathcal{L}_{H}$ governs the Hermitian dynamics, $\mathcal{L}_{D}$ represents the dissipative part, and $\mathcal{L}$ acts as a superoperator on the operator space.
+The inner product can be defined as $(A|B):=\mathrm{Tr}[\rho_{\infty}A^{\dagger}B]$, which explicitly connects the operator space structure to the Krylov basis.
+Secondly, an appropriate basis is required to compute the aforementioned physical quantities.
+Following Bhattacharya:2022gbz , we adopt an orthonormal basis $\{\mathcal{O}_{0},\ldots,\mathcal{O}_{n},\ldots\}$ generated by the open-system Lindbladian according to the following procedure:
+$\rm span(\mathcal{O}_{0},...,\mathcal{O}_{n})=span(\mathcal{V}_{0},...,\mathcal{V}_{n}),$ (10)
+where $\mathcal{V}_{n}$ is the original Krylov basis, and this basis can be generated by the Arnoldi iteration.
+Here, we will recall the main idea which is started with an initial normalized vector $\mathcal{O}_{0}\propto\mathcal{O}$.
+Then, we could construct the subsequent Krylov basis vectors by repeatedly applying the operator and orthonormalizing at each step.
+$|\mathcal{U}_{k})=\mathcal{L}|\mathcal{O}_{k-1}).$ (11)
+From $j=0$ to n1n-1, one can perfom the following iterations: $1.~h_{j,k-1}=(\mathcal{O}_{j}|\mathcal{U}_{k})$, $2.~|\tilde{\mathcal{U}}_{k})=|\mathcal{U}_{k})-\sum_{j=0}^{k-1}h_{j,k-1}|\mathcal{O}_{j})$, $3.~h_{k,k-1}=\sqrt{(\tilde{U}_{k}|\tilde{\mathcal{U}}_{k})}$, it will stop as if $h_{k,k-1}=0$, otherwise, we could define
+$|\mathcal{O}_{k})=\frac{|\tilde{\mathcal{U}}_{k})}{h_{k,k-1}}.$ (12)
+Consequently, the Lindbladian can transform into o an upper Heissenberg form in the span Krylov basis (also called the Arnoldi basis)
+$\mathcal{L}=\begin{pmatrix}h_{0,0}&h_{0,1}&h_{0,2}&\cdots&\cdots&h_{0,n}\\ h_{1,0}&h_{1,1}&h_{1,2}&\cdots&\cdots&h_{1,n}\\ 0&h_{2,1}&h_{2,2}&h_{2,3}&\cdots&\cdots\\ \vdots&0&h_{3,2}&\cdots&\cdots&\cdots\\ 0&\vdots&0&\vdots&\ddots&h_{n-1,n}\\ 0&0&\vdots&0&h_{n,n-1}&h_{n,n}\end{pmatrix}$ (13)
+where $h_{m,n}=(\mathcal{O}_{m}|\mathcal{L}|\mathcal{O}_{n})$.
+When $\mathcal{L}$ is a Hermitian superoperator, the Arnoldi iteration will nicely recover into the Lanczos algorithm.
+Following our previous conventions Li:2024kfm , we express the generalized Lanczos algorithm as
 $\mathcal{L}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1}),$ (14)
-where the coefficients $c_{n}$ encode the open quantum system dynamics for a given Hamiltonian, while $b_{n}$ (the Lanczos coefficients) characterize the chaotic properties of the dynamical system.
-The orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$ spans the Krylov space, where it denotes the space of the two-mode squeezed state, and $\mathcal{L}$ denotes the Liouvillian superoperator – which can be represented in terms of creation and annihilation operators analogous to a Hamiltonian operator.
-Following our previous work [25, 39, 40], we identify that:
-  • The $c_{n}$ terms in (6) correspond to open system effects
-  • The $b_{n}$ terms represent closed system dynamics
+where the coefficients $c_{n}$ capture the open quantum system dynamics for a given Hamiltonian (corresponding to the diagonal elements $h_{n,n}$ in (13)), while the Lanczos coefficients $b_{n}$ (associated with the off-diagonal elements $h_{n,n-1}$ in (13)) characterize the chaotic behavior of the dynamical system.
+The orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$ spans the Krylov space defined by Eq.
+(10).
+The Liouvillian superoperator $\mathcal{L}$, which admits a representation in terms of creation and annihilation operators equivalent to that of a Hamiltonian operator, which governs the time evolution.
 This formulation allows us to interpret $\mathcal{L}$ as an effective Hamiltonian that acts on two-mode quantum states, enabling explicit separation between open and closed system contributions.
@@
   1. 1.
-    Isolate environmental decoherence effects (open system)
+    Isolate the part of open system
   2. 2.
     Preserve unitary evolution (closed system)
-For comprehensive discussions of the Lanczos algorithm implementation and Krylov complexity in this context, we refer readers to [46].
-[…]
+For comprehensive discussions of the Lanczos algorithm implementation and Krylov complexity in this context, we refer readers to Parker:2018yvk .
+To connect with the generalized Lanczos algorithm, we need to represent the action (3) into the Hamiltonian in terms of creation and annihilation operators, where the definition of Hamiltonian is $H=\int d^{3}xd\eta(\pi v^{\prime}-L)$ with the conjugate momentum as
 $\pi(\eta,\vec{x})=\frac{\delta L}{\delta v^{\prime}(\eta,\vec{x})}=v^{\prime}-\frac{a^{\prime}}{a}v.$ (15)
 Therefore, we could obtain the Hamiltonian
-[…]
+$H=\frac{1}{2}\int d^{3}xd\eta[\pi^{2}+(\partial_{i}v)^{2}+\frac{a^{\prime}}{a}(\pi v+v\pi)].$ (16)
 Utilizing the Fourier decomposition for $v$ and $\pi$,
@@
 $\hat{\pi}(\eta,\vec{x})=i\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{k}{2}}(\hat{c}_{-\vec{k}}^{\dagger}u_{\vec{k}}^{\ast}e^{-i\vec{k\cdot}\vec{x}}-\hat{c}_{\vec{k}}u_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}).$ (18)
-where $\hat{c}_{-\vec{k}}^{\dagger}$ and $\hat{c}_{\vec{k}}$ denote the creation and annihilation operators, respectively, $u_{k}$ and $v_{k}$ are the function of conformal time $\eta$.
-Here, we keep two modes $(k,-k)$ since it is invariant under changing from $k$ to $-k$ in action (3) like in [36].
-Therefore, the Hamiltonian (8) will become as follows
-$\begin{split}\hat{H}=\int{d^{3}k}\hat{H}_{k}=&\int{d^{3}k}\bigg{[}k\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+k\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\\ &+i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}-i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}\bigg{]}.\end{split}$ (11)
-where have chosen appropriate normalizations for $v_{\vec{k}}$ and $u_{\vec{k}}$ and this Hamiltonian is of the Weyl-group structure [57].
-Then, we can use this Hamiltonian (11) to act on the two-mode orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$, one can obtain
-$\begin{split}\hat{H}|\mathcal{O}_{n})=(2n+1)k|\mathcal{O}_{n})+i\frac{z{}^{\prime}}{z}(n+1)|\mathcal{O}_{n+1})\\ -in\big{(}\frac{z{}^{\prime}}{z}\big{)}|\mathcal{O}_{n-1}).\end{split}$ (12)
-where $b_{n}^{2}=n^{2}(\frac{z^{\prime}}{z})^{2}$, $c_{n}=i(2n+1)k$ and the prime symbol indicates differentiation with respect to conformal time.
-The formula of Lanczos coefficient clearly indicates that the early universe is a maximally chaotic, infinity dynamical system as discovered in [25, 39, 40, 26].
-For convenience, we will define
+where $\hat{c}_{-\vec{k}}^{\dagger}$ and $\hat{c}_{\vec{k}}$ represent the creation and annihilation operators, respectively, and $u_{k}(\eta)$, $v_{k}(\eta)$ are functions of conformal time $\eta$.
+We retain both modes $(\vec{k},-\vec{k})$ due to the invariance of the action (3) under $\vec{k}\rightarrow-\vec{k}$, following the treatment in Agullo:2022ttg .
+Accordingly, the Hamiltonian (16) takes the following form:
+$\begin{split}\hat{H}=\int{d^{3}k}\hat{H}_{k}=\int{d^{3}k}\bigg{[}k\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+k\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}-i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}\bigg{]}.\end{split}$ (19)
+where we have chosen appropriate normalizations for $v_{\vec{k}}$ and $u_{\vec{k}}$ and this Hamiltonian is of the $SL(2,R)$ structure Caputa:2021sib .
+We then act with this Hamiltonian (19) on the two-mode orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$, yielding:
+$\begin{split}\hat{H}|\mathcal{O}_{n})=(2n+1)k|\mathcal{O}_{n})+i\frac{z{}^{\prime}}{z}(n+1)|\mathcal{O}_{n+1})-in\big{(}\frac{z{}^{\prime}}{z}\big{)}|\mathcal{O}_{n-1}).\end{split}$ (20)
+where
+$b_{n}^{2}=n^{2}(\frac{z^{\prime}}{z})^{2},~~~~c_{n}=i(2n+1)k,$ (21)
+The prime symbol denotes differentiation with respect to conformal time.
+As shown in Parker:2018yvk , the dynamical system becomes maximally chaotic and infinite when $b_{n}\propto n$.
+Hence, the expression for the Lanczos coefficient in (21) clearly demonstrates that the early universe constitutes a maximally chaotic, infinite dynamical system, as established in Li:2024kfm ; Li:2024iji ; Li:2024ljz ; Zhai:2024odw .
+For convenience, we define
 $M_{P}^{2}|1-\mu_{1}^{2}|=\bigg{(}\frac{z^{\prime}}{z}\bigg{)}^{2},~~M_{P}\mu_{2}=k,$ (22)
-where we have followed the notation of [25] and restored the unit of Planck mass, $\mu_{2}$ plays a role of dissipative coefficient and thus $b_{n}^{2}=|1-\mu_{1}^{2}|$ and $c_{n}=i(2n+1)\mu_{2}$.
-As previously mentioned, the $c_{n}$ part corresponds to the open system part of Hamiltonian (11), and the $b_{n}$ matches the closed system part of Hamiltonian (11).
-Thus, we could split the Hamiltonian into the open part and closed part as follows,
+where we follow the notation of Li:2024kfm and restore the Planck mass units.
+Here, $\mu_{2}$ acts as a dissipative coefficient, leading to the expressions $b_{n}^{2}=n^{2}|1-\mu_{1}^{2}|$ and $c_{n}=i(2n+1)\mu_{2}$.
+As noted earlier, the $c_{n}$ term corresponds to the open system contribution of the Hamiltonian (19), while the $b_{n}$ term aligns with the closed system part.
+We can therefore decompose the Hamiltonian into open and closed parts as follows:
 $\begin{split}&H_{open}=M_{P}\mu_{2}(\hat{c}_{-k}^{\dagger}\hat{c}_{-k}+\hat{c}_{k}\hat{c}_{k}^{\dagger}),\\ &H_{closed}=iM_{P}\sqrt{|1-\mu_{1}^{2}|}(\hat{c}_{k}^{\dagger}\hat{c}_{-k}^{\dagger}-\hat{c}_{k}\hat{c}_{-k}).\end{split}$ (23)
 For any single field inflationary models, its Hamiltonian can be divided into two parts as shown in Eq. (23), where the information of various models are encoded in $\mu_{2}$ and $\mu_{1}$.
+Here, we propose a new perspective on the generalized Lanczos algorithm in this work.
+This algorithm explicitly highlights the open system contribution to the quadratic perturbation action in (19), an aspect often overlooked in conventional treatments, where the dissipative coefficient is defined in direct relation to various Hamiltonian components.
 IV wave function
-As Ref.
-[47] pointed out, the amplification from the quantum level to the classical level in inflation is a quantum squeezing process.
-Therefore, it is natural to use the two-mode squeezed state to evaluate the quantum fluctuations in inflation.
-However, in this work, we will implement a more generic function to replace this wave function.
-First and foremost, the so-called OTMSS via Ref.
-[25] is
+As noted in Ref.
+Grishchuk:1990bj , the amplification from quantum to classical scales during inflation is fundamentally a quantum squeezing process.
+It is therefore natural to employ the two-mode squeezed state to describe quantum fluctuations in this context.
+In this work, however, we adopt a more general wave function.
+Specifically, we use the OTMSS, as introduced in Li:2024kfm , defined as follows:
 $|\mathcal{O}(\eta))=e^{i\mathcal{L}\eta}|0,0\rangle_{-\vec{k},\vec{k}}=\frac{\rm sechr_{k}(\eta)}{1+\mu_{2}\tanh r_{k}(\eta)}\sum_{n=0}^{\infty}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(-\exp(2i\phi_{k}(\eta))\tanh r_{k}(\eta))^{n}}{(1+\mu_{2}\tanh r_{k}(\eta))^{n}}|n,n\rangle_{-\vec{k},\vec{k}},$ (24)
-where $|n,n\rangle_{-\vec{k},\vec{k}}$ is the Krylov basis in this work, $\eta$ is the conformal time, $r_{k}$ and $\phi_{k}$ are two squeezing parameters, $\mu_{2}$ and $|1-\mu_{1}^{2}|$ as defined in Eq.
+where $|n,n\rangle_{-\vec{k},\vec{k}}$ denotes the Arnoldi basis used in this work, $\eta$ is the conformal time, $r_{k}$ and $\phi_{k}$ are the squeezing parameters, and $\mu_{2}$ and $|1-\mu_{1}^{2}|$ are as defined in Eq.
 (22).
-Here, we will introduce more information on wave function (15), in which this wave function will describe the quantum fluctuations in inflation.
-First, this wave function could represent the dissipative effects of one dynamical system denoted by $\mu_{2}$, and it will nicely recover the two-mode squeezed state in the weak dissipative approximation $(\mu_{2}\ll 1)$,
+We now provide further details on the wave function given in Eq.
+(24), which describes quantum fluctuations during inflation.
+This wave function incorporates dissipative effects through the parameter $\mu_{2}$, and reduces to the standard two-mode squeezed state in the weak dissipative limit ($\mu_{2}\ll 1$),
 $|\mathcal{O}(\eta))=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}+\mathcal{O}^{n}(\mu_{2}).$ (25)
-Thus, we can see that the pure two-mode squeezed state is a wave function that corresponds to a closed system.
-In later illustrations, we will confirm this point.
-Secondly, the derivation of the OTMSS (15) only relies on the second kind of Meixner polynomial [48],
+It follows that the pure two-mode squeezed state corresponds to a wave function describing a closed quantum system, a point we will reaffirm in subsequent discussions.
+Secondly, the OTMSS given in Eq.
+(24) is constructed based on the second kind of Meixner polynomial, Hetyei_2009 ,
 $P_{n+1}(x)=(x-\tilde{c}_{n})P_{n}(x)-b_{n}^{2}P_{n-1}(x),$ (26)
-where we define $\tilde{c}_{n}=-ic_{n}$, and $x$ denotew the Hamiltonian.
+where we define $\tilde{c}_{n}=-ic_{n}$, and $x$ denote the Hamiltonian.
 The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$.
@@
 $b_{n+1}e_{n+1}+b_{n}e_{n-1}=(x-\tilde{c}_{n})e_{n},$ (27)
-where $e_{n}=\mathcal{O}_{n}$ in our case and then it is equivalent to the generalized Lanczos algorithm (6).
-Through this simple illustration, one could explicitly see that the OTMSS only relies on the second kind of Meixner polynomial, describing the dissipative effects of open system (the early universe).
-Therefore, we could see that the OTMSS (15) is more generic compared with the pure two-mode squeezed state.
-Finally, we need to emphasize that the OTMSS is model-independent since the information of various models is encoded in $\mu_{2}$ and $|1-\mu_{1}^{2}|$.
+where $e_{n}=\mathcal{O}_{n}$ in our construction, which is equivalent to the generalized Lanczos algorithm given in Eq.
+(14).
+From this straightforward representation, it is evident that the OTMSS is constructed entirely from the second kind of Meixner polynomial, capturing dissipative effects inherent to open quantum systems such as the early universe.
+This demonstrates that the OTMSS in Eq.
+(24) is more general than the pure two-mode squeezed state.
+Moreover, the OTMSS is model-independent, as the parameters $\mu_{2}$ and $|1-\mu_{1}^{2}|$ encode specific physical information from different cosmological scenarios.
+In this section, we have already checked that the OTMSS is more generic and complete compared to the two-mode squeezed state.
+Before presenting the conceptual advantage of the OTMSS, we need to provide additional explanation for the two-mode squeezed state.
+To proceed, we will define a generalized displacement operator as follows:
+$\begin{split}D(\eta)&\ =e^{iHt}=\exp(ikt(c_{k}c_{k}^{\dagger}+c_{-k}^{\dagger}c_{-k})-\frac{z^{\prime}}{z}t(c_{-k}^{\dagger}c_{k}^{\dagger}-c_{-k}c_{k}))\\ &\ =\exp(-\eta c_{k}^{\dagger}c_{-k}^{\dagger})\exp(\bar{\eta}c_{-k}c_{k})\exp(-\alpha(c_{k}c_{k}^{\dagger}+c_{-k}^{\dagger}c_{-k})),\end{split}$ (28)
+where $t$ is just a parameter, $\hat{H}$ is our Hamiltonian operator (19) and we also define $\eta=\frac{z^{\prime}}{z}t-i\frac{z^{\prime}}{z}kt^{2}$ and $\alpha=\frac{1}{2}(\frac{z^{\prime}}{z}t-\frac{z^{\prime}}{z}kt^{2})^{2}-ikt$.
+After some tedious calculations, we could obtain
+$\begin{split}|\mathcal{O}(t))&\ =\exp(-\eta c_{k}^{\dagger}c_{-k}^{\dagger})\exp(\bar{\eta}c_{-k}c_{k})\exp(-\alpha(c_{k}c_{k}^{\dagger}+c_{-k}^{\dagger}c_{-k}))|0,0\rangle_{k,-k}\\ &\ =b\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}e^{i2n\phi}\tanh^{n}r(c_{-k}^{\dagger}c_{k}^{\dagger})^{n}|0,0\rangle_{k,-k}\\ &\ =b\sum_{n=0}^{\infty}(-1)^{n}e^{i2n\phi}\tanh^{n}r|n,n\rangle_{k,-k},\end{split}$ (29)
+where we have set $\eta=e^{2i\phi}\tanh r$ in complex coordinates, $r=|\eta|$ and $b$ is the normalization factor that can be derived as
+$1=(\mathcal{O}(t)|\mathcal{O}(t))=|b|^{2}\sum_{n=0}^{\infty}\tanh^{2n}r=|b|^{2}\frac{1}{1-\tanh^{2}r}=|b|^{2}\cosh^{2}r,$ (30)
+we obtain $b=\frac{1}{\cosh r}$.
+This result demonstrates that only the $H_{\text{closed}}$ component constructs the conventional two-mode squeezed state.
+Consequently, the OTMSS, derived from the second kind of Meixner polynomial, represents a generalization of the standard two-mode squeezed state.
+To our knowledge, this is the first time the OTMSS has been applied to compute the inflationary power spectrum.
+This generalization naturally incorporates dissipative effects, as described by the $H_{\text{open}}$ term in Eq.
+(23).
 V Bogoliubov transformation
-Since the power spectrum of BD vacuum is well consistent with observations [6], we will use the Bogoliubov transformation to construct the relation between the BD vacuum and the vacuum of (15).
-First, the mode function $v_{z}(\eta,\vec{k})$ parametrizes the squeezed vacuum state in time-evolving spacetime.
-As for the reference BD vacuum mode, it is denoted by $v_{BD}(\eta,\vec{k})$.
-Then, we could introduce the non-trivial Bogoliubov transformation to relate $v_{z}(\eta,\vec{k})$ and $v_{BD}(\eta,\vec{k})$ as follows,
+Since the power spectrum of the BD vacuum is in excellent agreement with observational data Planck:2018jri , we employ the Bogoliubov transformation to establish a relation between the BD vacuum and the vacuum state defined in Eq.
+(24).
+Specifically, the mode function $v_{z}(\eta,\vec{k})$ parameterizes the squeezed vacuum state in a time-dependent cosmological background.
+The reference mode function associated with the BD vacuum is denoted by $v_{\text{BD}}(\eta,\vec{k})$.
+We then introduce a non-trivial Bogoliubov transformation relating $v_{z}(\eta,\vec{k})$ and $v_{\text{BD}}(\eta,\vec{k})$ as follows:
 $v_{z}(\eta,\vec{k})=\alpha_{k}v_{BD}(\eta,\vec{k})+\beta_{k}v_{BD}^{*}(\eta,\vec{k}),$ (31)
@@
 $\hat{c}_{\vec{k}}=\alpha_{k}\hat{a}_{\vec{k}}+\beta_{k}^{*}\hat{a}_{-\vec{k}}^{\dagger}.$ (33)
-Then, the we could use the Bogoliubov coefficient to construct the wave function via BD vacuum as follows,
+Then, we could use the Bogoliubov coefficient to construct the wave function via BD vacuum as follows,
 $|0,0\rangle_{OST}=\frac{1}{\mathcal{N}}\exp\bigg{[}-\int\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\frac{\beta_{k}^{*}}{\alpha_{k}^{*}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\bigg{]}|BD\rangle,$ (34)
-where $\mathcal{N}$ represents the normalization factor, and $|0,0\rangle_{OST}$ denotes the vacuum of the OTMSS.
+where $\mathcal{N}$ is the normalization factor, and $|0,0\rangle_{\text{OST}}$ denotes the vacuum of the OTMSS, which is explicitly constructed from the orthogonal basis $|n,n\rangle_{-\vec{k},\vec{k}}$ generated via specific recurrence relations Gamayun:2025hvu ; Muck:2022xfc .
 The essence of Eq.
-(22) is that one can construct the wave function in the target spacetime using the Bogoliubov transformation and the BD vacuum.
-Therefore, the key aspect is determining the formulas for $\alpha_{k}$ and $\beta_{k}$ for our specific case.
-This is also crucial for understanding the relationship between the modes of the squeezed state and the modes of the BD vacuum, as indicated in Eq.
+(34) is that the wave function in the target spacetime can be systematically constructed using the Bogoliubov transformation applied to the BD vacuum.
+It is therefore essential to determine the explicit form of the Bogoliubov coefficients $\alpha_{k}$ and $\beta_{k}$ in our framework.
+This is crucial for understanding the correspondence between the squeezed state modes and those of the BD vacuum, as articulated in Eq.
 (35).
-In fact, we can apply the squeezed operator $\mathcal{\hat{S}}_{-k,k}=\exp\left(\frac{r_{k}}{2}e^{-i\phi_{k}}\hat{a}_{k}\hat{a}_{-k}-\frac{r_{k}}{2}e^{i\phi_{k}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$ to the annihilation operator of the BD vacuum to derive the coefficients for the Bogoliubov transformation,
+In practice, we apply the two-mode squeezing operator
+$\mathcal{\hat{S}}_{-k,k}=\exp\left(\frac{r_{k}}{2}e^{-i\phi_{k}}\hat{a}_{k}\hat{a}_{-k}-\frac{r_{k}}{2}e^{i\phi_{k}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$
+to the annihilation operators of the BD vacuum to derive the Bogoliubov coefficients.
 $\mathcal{\hat{S}}^{\dagger}\hat{a}\mathcal{\hat{S}}=\hat{c}_{\vec{k}}=\cosh r_{k}\hat{a}_{\vec{k}}-e^{i\phi_{k}}\sinh r_{k}\hat{a}_{-\vec{k}}^{\dagger},$ (35)
@@
 $\alpha_{k}=\cosh r_{k},~~~\beta_{k}=-\exp(-i\phi_{k})\sinh r_{k},$ (36)
-where $\phi_{k}$ and $r_{k}$ are two parameters depicting the properties for the OTMSS (15) and the formula of $\alpha_{k}$ and $\beta_{k}$ is consistent with [49, 50].
+where $\phi_{k}$ and $r_{k}$ are two parameters depicting the properties for the OTMSS (24) and the formula of $\alpha_{k}$ and $\beta_{k}$ is consistent with Bianchi:2024jmn ; Grain:2019vnq .
 Next, we will proceed with the evolution of $\phi_{k}$ and $r_{k}$ determined by the Hamiltonian (19).
-As we discussed, the OTMSS (15) is more generic compared with a pure two-mode squeezed state, including the dissipative effects of an open system.
-In this work, we will particularly focus on the super Hubble scale, namely, $k\eta$ will be approaching zero.
-In light of (13), we could see that $\mu_{2}$ (dissipative coefficient) will also be very tiny in this super Hubble scale, meaning that the universe within the super Hubble scale is a weak dissipative system.
-Consequently, the pure two-mode squeezed state will be a good approximation, and our derivation for (24) is also consistent with each other.
+As previously established, the OTMSS in Eq.
+(24) represents a generalization of the pure two-mode squeezed state, as it incorporates dissipative effects inherent to open quantum systems.
+In this work, we focus specifically on the super-Hubble regime, where $k\eta$ approaches zero.
+According to (22), $\mu_{2}$ (the dissipative coefficient) becomes extremely small in this regime, indicating that the universe at the super-Hubble scale behaves as a weakly dissipative system.
+Therefore, the pure two-mode squeezed state serves as a good approximation, and our derivation for (36) remains consistent with this understanding.
+In the sub-horizon regime, as $k\eta$ approaches infinity, the dissipative coefficient $\mu_{2}$ becomes the unity as discovered by Zhai:2024odw ; Li:2024kfm ; Li:2024iji ; Li:2024ljz This indicates that the sub-horizon region acts as a strongly dissipative system, leading to that we should use the OTMSS to investigate the observable by taking place of the two-mode squeezed state.
+Consequently, we expect observable predictions to deviate from those of the standard scenario.
+This provides a physical interpretation of observational consequences throughout the entire inflationary epoch: initially, observable outcomes deviate from the standard case in the sub-horizon region.
+After horizon exit, when the mode enters the super-horizon regime, the observational signatures align with the standard results.
+In the following calculations, we will use the power spectrum to illustrate this point in the super-horizon regime since it is related to the observations.
 VI Evolution of $\phi_{k}$ and $r_{k}$
-To obtain the power spectrum of the OTMSS (15), one needs to rewrite the Bogoliubov coefficients (24) in terms of $k$ as in Refs.
-[51, 52, 53], where their formulas has made some approximation.
-To fully capture the information of (15), we will explicitly numerically give the evolution of $r_{k}$ and $\phi_{k}$ in terms of $k$, where the evolution is completely determined by the Schrödinger equation $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$.
-The resulting evolution equation of $r_{k}$ and $\phi_{k}$ can be denoted by via [26]
+To derive the power spectrum of the OTMSS given in Eq.
+(24), the Bogoliubov coefficients in Eq.
+(36) must be expressed as functions of the wavenumber $k$, following the approach adopted in Refs.
+Danielsson:2002kx ; Danielsson:2004xw ; Broy:2016zik , where certain approximations were employed.
+To fully incorporate the structure of the OTMSS wave function (24), we numerically compute the evolution of the squeezing parameters $r_{k}$ and $\phi_{k}$ with respect to $k$.
+This evolution is governed by the Schrödinger equation
+$\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle,$
+which completely determines the dynamics. The resulting equations of motion for $r_{k}$ and $\phi_{k}$ can be expressed as derived in Zhai:2024odw :
 $r_{k}^{\prime}=\frac{-M_{P}|1-\mu_{1}^{2}|\sinh 2r_{k}\cos 2\phi_{k}-\sinh 2r_{k}\cdot\mu_{2}^{\prime}}{\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k}},$ (37)
 $\phi_{k}^{\prime}=-M_{P}\mu_{2}+\frac{1}{2}M_{P}|1-\mu_{1}^{2}|^{\frac{1}{2}}\sin 2\phi_{k}\Big{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\coth r_{k}+\mu_{2})\Bigr{]}.$ (38)
-Compared to the original evolution equations for $\phi_{k}$ and $r_{k}$, we have neglected the contribution from the potential due to the slow-roll conditions.
-These calculations can be quite complicated; for more details, the reader can refer to the appendix of [26].
-The evolution equations (26) for $r_{k}$ and $\phi_{k}$ describe the dynamics during the inflationary period.
-The power spectrum was observed at the horizon crossing at a fiducial momentum of $k_{*}=0.05~{}\text{Mpc}^{-1}$.
-To compare with observations, the data from the CMB covers the range from $10^{-4}~{}\text{Mpc}^{-1}$ to $1~{}\text{Mpc}^{-1}$ [6], making it beneficial to fully expand the power spectrum within this range.
-At the pivot point, we have $k_{*}=aH$, where the value is derived from today’s measurements, allowing us to easily obtain $\eta=-\frac{1}{k}$ within the CMB observational range [54, 55, 56].
-Given that $\eta=-\frac{1}{aH}$, the evolution can be transformed accordingly,
-$\begin{split}r_{k}^{\prime}=&\ -\frac{\tanh r_{k}(\mu_{2}^{\prime}+M_{P}|1-\mu_{1}^{2}|\cos 2\phi_{k})}{\tanh r_{k}+\mu_{2}},\\ \phi_{k}^{\prime}=&\ -M_{P}\mu_{2}+\frac{1}{2}\sin 2\phi_{k}\Bigl{(}\frac{M_{P}|1-\mu_{1}^{2}|\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\\ &\ +M_{P}\coth r_{k}+M_{P}\mu_{2}\Bigr{)}.\end{split}$ (27)
-Being armed with these two equations, we could give their numeric in terms of comoving momentum $k$.
+Compared to the original evolution equations for $\phi_{k}$ and $r_{k}$, we have neglected the contribution from the potential term due to the slow-roll conditions.
+These calculations can be rather involved; for complete details, we refer the reader to the appendix of Zhai:2024odw .
+The evolution equations (38) for $r_{k}$ and $\phi_{k}$ describe the squeezing dynamics during the inflationary epoch.
+The power spectrum is evaluated at horizon crossing for a fiducial scale $k_{*}=0.05~\text{Mpc}^{-1}$.
+To facilitate comparison with observations, we note that CMB measurements cover the range from $10^{-4}~\text{Mpc}^{-1}$ to $1~\text{Mpc}^{-1}$ Planck:2018jri .
+It is therefore useful to compute the power spectrum across this full range.
+Under these considerations, Eq.
+(38) becomes:
+$\begin{split}r_{k}^{\prime}=&\ -\frac{\tanh r_{k}(M_{P}|1-\mu_{1}^{2}|\cos 2\phi_{k})}{\tanh r_{k}+\mu_{2}},\\ \phi_{k}^{\prime}=&\ -M_{P}\mu_{2}+\frac{1}{2}\sin 2\phi_{k}\Bigl{(}\frac{M_{P}|1-\mu_{1}^{2}|\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+M_{P}\coth r_{k}+M_{P}\mu_{2}\Bigr{)}.\end{split}$ (39)
+To obtain the corresponding numerical values of $r_{k}$ and $\phi_{k}$, we use the horizon-crossing condition $k_{*}=aH$ at the pivot scale, whose value is inferred from present-day observations.
+This allows us to express the conformal time as $\eta=-1/k$ across the CMB observational range Bianchi:2024qyp ; Gong:2001he ; Auclair:2022yxs .
+Using these evolution equations, we numerically compute $r_{k}$ and $\phi_{k}$ as functions of the comoving momentum $k$.
 Figure 1: The numeric of squeezed parameter $r_{k}$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
@@
 3 clearly indicates that mean number density is almost zero since the particles are generated in preheating process, although the number density of particle is non-vanishing.
-Being armed with the evolution of $r_{k}$ and $\phi_{k}$ in terms of k, one can investigate the power spectrum.
+Using the evolution of $r_{k}$ and $\phi_{k}$ in terms of k, one can investigate the power spectrum.
 VII Power spectrum of OTMSS
-Since the power spectrum of OTMSS (15) is proportional to $|v_{z}|^{2}$. Then, we can obtain the ratio between the power spectrum of the OTMSS and the BD vacuum in light of (23),
+Since the power spectrum of the OTMSS in Eq.
+(24) scales with $|v_{z}|^{2}$, we may express the ratio between this power spectrum and that of the BD vacuum as follows, using the Bogoliubov transformation given in Eq.
+(35):
 $\gamma_{z}=\frac{\Delta_{\mathcal{R}z}^{2}}{\Delta_{\mathcal{R}}^{2}}=|\alpha-\beta|^{2}=\cosh 2r_{k}+\sinh 2r_{k}\cos\phi_{k},$ (41)
-where $\Delta_{\mathcal{R}z}^{2}$ denotes the power spectrum for the OTMSS, $\Delta_{\mathcal{R}}^{2}$ represents the power spectrum of BD vacuum and we have used $\lim_{k\eta\to 0}\frac{v_{BD}^{*}}{v_{BD}}=-1$ (super Hubble scale approximation) via BD vacuum (5).
-Once having this relation, it is quite straightforwardly showing the numeric of (29) in terms of k.
+where $\Delta_{\mathcal{R}z}^{2}$ denotes the power spectrum for the OTMSS, $\Delta_{\mathcal{R}}^{2}$ represents the power spectrum of the BD vacuum, and we have used the super-Hubble scale approximation $\lim_{k\eta\to 0}v_{BD}^{*}/v_{BD}=-1$ derived from the BD vacuum definition in Eq.
+(5).
+With this relation established, it is straightforward to compute the numerical evaluation of Eq.
+(41) as a function of $k$.
 Figure 4: The numeric of squeezed parameter $\gamma_{z}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
 Fig.
-4 indicates that the ratio (29) is almost one after the horizon crossing, inferring that the power spectrum of two-mode squeezed state will be consistent with the BD vacuum.
-For completeness, we will use the definition of power to confirm this conclusion.
-Following the notation of [34], the two-point correlation function can be defined as
+4 shows that the ratio in Eq.
+(41) is very close to unity after horizon crossing, which implies that the power spectrum of the two-mode squeezed state is consistent with that of the BD vacuum.
+For completeness, we further validate this result using the formal definition of the power spectrum.
+Following the notation of Baumann:2009ds , the two-point correlation function is defined as
 $\begin{split}\langle|\hat{v}_{z}|^{2}\rangle=&\ \langle 0,0|\hat{v}_{z}^{\dagger}(\eta,0)\hat{v}_{z}(\eta,0)|0,0\rangle\\ =&\ \int d\ln{k}\frac{k^{3}}{2\pi^{2}}|v_{BD}(\eta,\vec{k})|^{2}\gamma_{z}.\end{split}$ (42)
@@
 $\begin{split}\Delta_{v_{z}}^{2}(\eta,k)=&\ \frac{k^{3}}{2\pi^{2}}|v_{BD}(\eta,\vec{k})|^{2}\gamma_{z}\\ =&\ \Delta_{v_{BD}}^{2}(\eta,k)\gamma_{z},\end{split}$ (43)
-To match the observations, we will use the curvature perturbation to calculate the power spectrum, where the curvature perturbation is $\mathcal{R}=v/z$ with $z=\sqrt{2\epsilon}a$ as shown in Sec.
+To compare with observational data, we compute the power spectrum using the curvature perturbation $\mathcal{R}=v/z$, where $z=\sqrt{2\epsilon}\,a$ as defined in Sec.
 II.
-Thus, we could obtain the power spectrum of curvature perturbation as follows,
+The power spectrum of the curvature perturbation is therefore given by:
 $\Delta_{\mathcal{R}z}^{2}(\eta,k)=\frac{1}{2\epsilon a^{2}}\frac{\Delta_{v_{z}}^{2}(\eta,k)}{M_{P}^{2}}=\frac{1}{2\epsilon a^{2}}\frac{\Delta_{v_{BD}}^{2}\gamma_{z}(\eta,k)}{M_{P}^{2}},$ (44)
-As for the power spectrum of BD vacuum, it will predict the $\Delta_{\mathcal{R}}^{2}=A_{s}(\frac{k}{k_{*}})^{n_{s}-1}$ where the observational constraints to the amplitude power spectrum $A_{s}$ and spectral index from [6]
+As for the power spectrum of the BD vacuum, it takes the form
+$\Delta_{\mathcal{R}}^{2}=A_{s}\left(\frac{k}{k_{*}}\right)^{n_{s}-1},$
+where the amplitude $A_{s}$ and the spectral index $n_{s}$ are constrained by observational data from Planck:2018jri .
 $A_{s}=(2.196\pm 0.060)\times 10^{-9},~~~n_{s}=0.9649\pm 0.0042.$ (45)
 Based on these observational constraints, we could give the prediction for the power spectrum of OTMSS in light of (44).
-Figure 5: The numeric of squeezed parameter $\Delta_{\mathcal{R}z}^{2}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
+Figure 5: The numeric of power spectrum $\Delta_{\mathcal{R}z}^{2}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
 Fig.
-5 clearly shows the power spectrum of the OTMSS is almost the same with BD vacuum due to the ratio relation (29) as shown in Fig.
-4, meanwhile our evaluation for the spectral index $n_{s}$ is also consistent with the BD vacuum.
-Recalling the calculation of the power spectrum for the squeezed state as outlined in [51, 52, 53], we can parametrize the Bogoliubov coefficients, $\alpha_{k}$ and $\beta_{k}$, in terms of $r_{k}$ and $\phi_{k}$.
-These coefficients will be expressed as functions of the momentum $k$.
-However, these procedures overlook a crucial aspect: both $r_{k}$ and $\phi_{k}$ evolve over time as determined by the corresponding Schrödinger equation, given by $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$.
-Therefore, a more accurate approach would be to take the time evolution of $r_{k}$ and $\phi_{k}$ into consideration.
-Furthermore, we focus on the OTMSS, as described in (15), which represents the open system (the universe is regarded as an open system).
-Following this approach, our numerical results for the power spectrum are consistent with the BD vacuum.
-While our derivation cannot be distinguished from the BD vacuum, the wave function in (15) is more general, being model-independent and relying solely on the second Meixner polynomial and the Lanczos algorithm.
+5 clearly demonstrates that the power spectrum of the OTMSS is nearly identical to that of the BD vacuum, which is consistent with the near-unity ratio shown in Eq.
+(41) and Fig.
+4.
+Furthermore, our evaluation of the spectral index $n_{s}$ agrees with the BD vacuum results within observational uncertainties.
+Recalling the calculation of the power spectrum for the squeezed state as outlined in Danielsson:2002kx ; Danielsson:2004xw ; Broy:2016zik , we may parametrize the Bogoliubov coefficients $\alpha_{k}$ and $\beta_{k}$ in terms of the squeezing parameters $r_{k}$ and $\phi_{k}$, which are in turn expressed as functions of the momentum $k$.
+However, these treatments neglect an essential feature: both $r_{k}$ and $\phi_{k}$ evolve in time according to the Schrödinger equation.
+A more complete approach must therefore incorporate this dynamical evolution.
+Moreover, we focus on the OTMSS defined in Eq.
+(24), which inherently describes an open quantum system—reflecting the fact that the universe is best treated as an open system.
+Within this framework, our numerical results for the power spectrum align with those of the BD vacuum.
+Although our predictions are observationally indistinguishable from the BD vacuum, the wave function in Eq.
+(24) is more general: it is model-independent and constructed solely from the second kind of Meixner polynomial within the generalized Lanczos algorithm.
 VIII Conclusion and Discussion
-Under the single field inflation, we have constructed a framework for calculating the power spectrum of two mode squeezed state, where the key ingredient is the OTMSS (15) and the generalized Lanczos algorithm as shown in Sec.
+Within the framework of single-field inflation, we have developed a method for computing the power spectrum of the two-mode squeezed state.
+The central elements of this construction are the OTMSS given in Eq.
+(24) and the generalized Lanczos algorithm, as detailed in Sec.
 III.
-In light of the Lanczos algorithm, the Hamiltonian can be separated into components corresponding to open and closed systems, as shown in Eq.
+Within the framework of the generalized Lanczos algorithm, the Hamiltonian can be decomposed into open and closed system components, as expressed in Eq.
 (23).
-During inflation, an amplification process drives quantum phenomena toward classicalization, making the two-mode squeezed state a natural framework for studying curvature perturbations.
-The squeezed state is constructed via the action of the squeezed operator $\mathcal{\hat{S}}_{-k,k}=\exp\left(u\hat{a}_{k}\hat{a}_{-k}-\bar{u}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$ on the BD vacuum, where $u=\frac{r_{k}}{2}e^{-i\phi_{k}}$ in our case.
-Crucially, (14) reveals that $\mathcal{\hat{S}}_{-k,k}$ exclusively encodes the closed system dynamics.
-In contrast, the OTMSS in (15) incorporates environmental interactions through the dissipative coefficient $\mu_{2}$.
-The standard squeezed state emerges as the leading-order term in the weak dissipation limit ($\mu_{2}\to 0$).
-From the Schrödinger equation $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$, we derive the evolution equations for $r_{k}(\eta)$ and $\phi_{k}(\eta)$, whose numerical solutions are plotted in Figs.
+During inflation, an amplification mechanism drives the quantum-to-classical transition, making the two-mode squeezed state a natural basis for characterizing curvature perturbations.
+This squeezed state is constructed by applying the squeezing operator
+$\mathcal{\hat{S}}_{-k,k}=\exp\left(u\hat{a}_{k}\hat{a}_{-k}-\bar{u}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$
+to the BD vacuum, where $u=\frac{r_{k}}{2}e^{-i\phi_{k}}$ in our notation.
+Crucially, Eq.
+(23) indicates that $\mathcal{\hat{S}}_{-k,k}$ captures only the closed system dynamics.
+In contrast, the OTMSS defined in Eq.
+(24) incorporates environmental interactions via the dissipative coefficient $\mu_{2}$.
+The conventional squeezed state is recovered in the weak dissipation limit ($\mu_{2}\to 0$) as the leading-order approximation.
+From the Schrödinger equation
+$\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle,$
+we derive the evolution equations for $r_{k}(\eta)$ and $\phi_{k}(\eta)$, with numerical solutions illustrated in Figs.
 2 and 1.
-The ratio between the OTMSS and BD vacuum, defined via Bogoliubov transformations in Eq.
-(29).
-As shown in Fig.
-4, $\gamma_{k}$ remains unity throughout inflation, consistent with suppressed particle production (mean particle density $\langle n_{k}\rangle\ll 1$) due to preheating dynamics.
-Consequently, the power spectrum $\Delta_{\mathcal{R}}^{2}(k)$ computed from the OTMSS in (15) matches the BD vacuum prediction, as demonstrated in Fig.
+The ratio $\gamma_{k}$ between the OTMSS and BD vacuum power spectra, defined via Bogoliubov transformations in Eq.
+(41), remains unity throughout inflation, as shown in Fig.
+4.
+This is consistent with strongly suppressed particle production (mean particle density $\langle n_{k}\rangle\ll 1$) due to preheating dynamics.
+Consequently, the power spectrum $\Delta_{\mathcal{R}}^{2}(k)$ computed from the OTMSS in Eq.
+(24) matches the prediction of the BD vacuum, as confirmed in Fig.
 5.
@@
   • Lanczos-based separation of $\hat{H}$ clarifies the open system’s part and closed system’s part.
-The derivation of the BD vacuum in (5) relies on solving the mode equation.
-In contrast, the OTMSS is constructed through the Hamiltonian formalism and Lanczos algorithm without solving equations of motion, where model-dependent information is encoded in the parameters $\mu_{2}$ and $|1-\mu_{1}^{2}|$.
-Remarkably, the power spectrum prediction from the OTMSS in (15) matches the BD vacuum result, suggesting that squeezed state construction is a more fundamental approach than solving EOM.
-The key ingredient in this construction is the generalized displacement operator [57], defined as
+The derivation of the BD vacuum in Eq.
+(5) relies on explicitly solving the mode equation of motion.
+In contrast, the OTMSS is systematically constructed within the Hamiltonian formalism via the Lanczos algorithm without solving the equations of motion; instead, model-dependent information is encoded in the parameters $\mu_{2}$ and $|1-\mu_{1}^{2}|$.
+Remarkably, the power spectrum predicted by the OTMSS in Eq.
+(24) matches the result from the BD vacuum, indicating that the squeezed-state formulation provides a more fundamental description than direct calculation of the equations of motion.
+The key ingredient in this construction is the generalized displacement operator Caputa:2021sib , defined as
 $\mathcal{D}(\xi)=\exp\left(\xi\hat{H}-\bar{\xi}\hat{H}^{\dagger}\right),$ (46)
-where $\hat{H}$ exhibits specific group structures (e.g., the Weyl algebra in our case) and the squeezed operator belongs this kind of operator.
+where $\hat{H}$ exhibits specific group structures (e.g., the $SL(2.R)$ in our case) and the squeezed operator belongs this kind of operator.
 When $\hat{H}$ belongs to a Lie group, the operator $\mathcal{D}(\xi)$ provides a rigorous framework for generating quantum states, surpassing traditional EOM-based methods in universality.
 Extensions and implications:
-  • Current results apply to single-field inflation. Generalizations to multifield scenarios [58, 59, 60, 61, 62] and $f(R)$ gravity [63, 64] are natural next steps.
+  • Current results apply to single-field inflation.
+Generalizations to multifield scenarios Liu:2019xhn ; Liu:2020zzv ; Liu:2020zlr ; Liu:2021rgq ; Zhang:2022bde and $f(R)$ gravity Liu:2018hno ; Liu:2018htf are natural next steps.
   • The $\mu_{2}$-dependence in OTMSS may encode decoherence effects observable in higher-order correlations.
+Pronounced dissipative effects on sub-horizon scales are expected to cause deviations in non-Gaussian correlators from those predicted by the standard two-mode squeezed state.
+These significant dissipative effects may also influence decoherence.
   • Group-theoretic construction offers a unified approach for quantum fluctuations in curved spacetime.
-Acknowledgements.
+Acknowledgements
 LHL and KHZ are funded by National Natural Science Foundation of China (NSFC) with grant NO.